OpenStudy (anonymous):

3x-4y=-9 -x-4y=-13 Solve these linear systems by graphing, then find the point at which they intersect. HOW DO I DO THIS!? i need help. -_-

OpenStudy (saifoo.khan):

Graph both lines.

OpenStudy (saifoo.khan):

they will definitely meet somewhere

OpenStudy (anonymous):

how do i graph them, what do i put on the x axis and what do i put on the y?

OpenStudy (saifoo.khan):

3x-4y=-9 graph this line first.

OpenStudy (saifoo.khan):

whn x-intercepts y = 0 whn y-intercpets x = 0

OpenStudy (saifoo.khan):

insert the values to get the intercepts.

OpenStudy (anonymous):

but then for the y intercept you get: 3x-4y=-9 3(0)-4y=-9 4y=-9 then you get -2.25, correct?

OpenStudy (anonymous):

dividing both sides by 4/

OpenStudy (saifoo.khan):

its correct.

OpenStudy (saifoo.khan):

now after y-intercpet. go for x intercpt & draw a line there.

OpenStudy (anonymous):

okay thank you so much mr.khan! lol

OpenStudy (saifoo.khan):

lol, uw.

OpenStudy (saifoo.khan):

Please hit good ans.